a freight train leaves the train station 1 hour before a passenger train. The two trains are traveling in the same direction on parallel tracks. if the rate of the passenger train is 5 mph than the freight train, how fast is each train traveling if the passenger train passes the freight train in 12 hours?
Answer by ankor@dixie-net.com(22740)

a freight train leaves the train station 1 hour before a passenger train.
The two trains are traveling in the same direction on parallel tracks.
if the rate of the passenger train is 5 mph more than the freight train,
how fast is each train traveling if the passenger train passes the freight train in 12 hours?
:
When the pass train catches the freight, they will have traveled the same dist.
From the information given we know the travel time of the pass is 12 hrs and the travel time of the freight is 13 hrs
:
let s = the speed of the freight
then
(s+5) = speed of the pass train
:
Write a distance equation: dist = speed * time
13s = 12(s+5)
13s = 12s + 60
13s - 12 = 60
s = 60 mph is the speed of the freight
You can find the speed of the pass
;
:
Pretty easy, right?
:
:
YOu can confirm this by finding the actual dist each train went, should be the same.